5^x=2e^(x+1) Solve for x
Answer by swincher4391(1107)

5^x = 2e^(x+1)
5^x = (2e^x)*(e)
(5^x)/(2e) = e^x
(1/2e) = e^x / 5^x
(1/2e) = (e/5)^x
(1/2e) = log(e/5)(e/5)^x
log(e/5)(1/2e) = x
Change of base formula
log(1/2e)/log(e/5) = x
x = 2.7782 <---- Answer